The length of a rectangle is 5 feet less than four times the width if the perimeter is 80 feet find the length and the width of the rectangle
Answer by algebra hello(55)

let width of rectangle=b
then length of rectangle(l)=4b-5
perimeter of rectangle=80
2(l+b)=80
2(4b-5+b)=80
2(5b-5)=80
10(b-1)=80
b-1=8
b=9
so width of rectangle(b)=9 feet